OpenStudy (callisto):

How to solve this equation? \[z^{z^{z^{\cdot ^{\cdot^{\cdot}}}}} = 2\] Update: By solving \(z^{z^{z^{\cdot ^{\cdot^{\cdot}}}}} = 2\) using the method we discussed in this post, we get \(z=\sqrt{2}\). Now, when we solve \(z^{z^{z^{\cdot ^{\cdot^{\cdot}}}}} = 4\) using the same method, we get \(z=\sqrt{2}\). So, the left side of the two equations are the same as \(z=\sqrt{2}\) in both equations. As a result, the right side should also be the same, which means we have 2=4. The problem is... what's wrong?

OpenStudy (callisto):

Please post "z^{z^{z^{\cdot ^{\cdot^{\cdot}}}}} = 2" in the LaTeX equation editor at http://www.codecogs.com/latex/eqneditor.php to read the equation. Thanks!

OpenStudy (anonymous):

can't get latex to work :C

OpenStudy (tkhunny):

One logarithm might suggest a solutions. \(z^{z^{z^{.^{.{.}}}}}\cdot z = log(2)\)

OpenStudy (anonymous):

wow

OpenStudy (tkhunny):

Whoops. That's log(z) on the left.

OpenStudy (solomonzelman):

Callistro, you want this? \[\huge\color{blue}{ z^{z^{z^z}} =2 }\]

OpenStudy (callisto):

@SolomonZelman No, not exactly. There are infinitely z's in the index part.

OpenStudy (tkhunny):

\(2\cdot \log(z) = \log(2)\) Almost done.

OpenStudy (tkhunny):

LaTeX is working fine form me. 2*log(z) = log(2)

OpenStudy (solomonzelman):

Sorry, got disconnected, do you know why latex disconnects me so much? \[\huge\color{blue}{ z^{z^{z^{.^{.^{.}}}}}=2 } \] sorry for the late reply.

OpenStudy (callisto):

Ah! So, here it is. \[z^{z^{z^{\cdot ^{\cdot^{\cdot}}}}} = 2\]\[z^{z^{z^{\cdot ^{\cdot^{\cdot}}}}}\log z = \log 2\]Replace \(z^{z^{z^{\cdot ^{\cdot^{\cdot}}}}}\) by 2 again in the equation, we get\[2\log z = \log 2\]\[\log z = \frac{1}{2} \log 2\]\[z = 2^{\frac{1}{2}}\] Is that right?

OpenStudy (tkhunny):

The arithmetic is good. Can you prove the sequence of exponents actually converges?

OpenStudy (callisto):

Would you mind giving me any hints to do it? :)

OpenStudy (tkhunny):

You cannot just keep raising a store number to the square root of 2. That is not the right structure. I'm feeling like L.Euler, today. Let's just assume infinite sequences work and move on. I did forget to ask. Is z a Real Number? Should we be looking for Complex Solutions?

OpenStudy (tkhunny):

10 Zs - 1.983668399 20 Zs - 1.999585623 40 Zs - 1.99999972847903 80 Zs - 1.99999999999988 88 Zs - 1.99999999999999 -- Ran out of that system's machine precision. If I had to vote, I would vote for convergence.

OpenStudy (callisto):

Just real number. My teacher said we are mostly deal with real number system in this course. Though, finding the solution to the equation is actually the end of the question.

OpenStudy (tkhunny):

This should strike wonder to the deepest part of your mathematical soul. An exponential tower of Irrational Numbers is a Rational Number???!!!!

OpenStudy (callisto):

And even worse, he used this to show 2=4...

OpenStudy (tkhunny):

Awesome. Maybe we were worrying about convergence from time to time.

OpenStudy (callisto):

Here is the problem he showed us in the first lecture: Solving \(z^{z^{z^{\cdot ^{\cdot^{\cdot}}}}} = 2\), we get \(z=\sqrt{2}\) Solving \(z^{z^{z^{\cdot ^{\cdot^{\cdot}}}}} = 4\), we get \(z=\sqrt{2}\) So, the left side of the two equations are the same as \(z=\sqrt{2}\) in both equations. As a result, the right side should also be the same, which means we have 2=4. The problem is... what's wrong? He said he hoped we could find out the problem at the end of the course, but I guess I can't wait that long.

OpenStudy (anonymous):

First z cannot be > 1, because the tower of power of z will go to infinity.

OpenStudy (anonymous):

If z =1, then the tower is on1

OpenStudy (anonymous):

If z <1, then this tower stay less than one. What is wrong with my thinking?

OpenStudy (callisto):

It... seems right to me...

OpenStudy (anonymous):

What was wrong in the reasoning to show that\( z=\sqrt 2 \) is that you use the fact that this infinite tower of powers of z converges. It does not always.

OpenStudy (anonymous):

@Callisto

OpenStudy (callisto):

I am sorry that I don't quite understand the step "ln(y_n)= z^n \ln(z)"

OpenStudy (experimentx):

I think this is due to singularity at x=1 ...

OpenStudy (anonymous):

Do it by induction

OpenStudy (callisto):

I am so confused... Actually, it goes back to the question, how do you know if this equation can be solved or not. What is wrong with previous solution to solve z? Why does taking log on both sides use the fact that the infinite tower of powers of z converges (which is not true, but I need time to figure out.)? I'm still trying to understand Dr's proof. Sorry for being so stupid.

OpenStudy (experimentx):

no no ... in analysis, you should prove that if you take infinitely many hyperpower ... you should show that it's limit exist before you find it's value. value come after you show that it exists. I don't understand either ... no latex here :(

OpenStudy (experimentx):

lol .... my grammar!!

OpenStudy (callisto):

One of the problems is that I haven't taken any analysis courses yet. This questions was raised in my Linear Algebra course.

OpenStudy (experimentx):

according to big bro ... the sequence of function only converges in the range .. (e^(-e), e^(1/e)) ...

OpenStudy (callisto):

I am not quite sure if I can understand the step "\(\ln(y_2)=z \ln (z^z)=z^2 \ln( z)\)" For \(y_2\), isn't it \(y_2 = z^{z^{z}}}\)? If it is, then wouldn't we get \(\ln(y_2)=z^z \ln (z)\) instead?

OpenStudy (anonymous):

You are right, I am wrong @Callisto

OpenStudy (experimentx):

let \[ f(x) = x^{x^{ .. }} = a \\ x^{x^{ .. }} \log(x) = \log(a) \\ a \log(x) = \log(a) \hspace{1 cm} -(1)\\ \log(x) = \frac 1 a \log(a) \] differentiating both sides, \[ \frac 1 x \frac{d x }{da} = \frac 1 {a^2 } (1 - \log(a)) \] upon maximizing it, we get \(a = e\) hence the max value of \( x \) is \( e^{1/e} \) .... beyond that value, \( f(x) \) does not converge. for \( x = \sqrt 2 \) , \( (1) \) obviously has two roots, http://www.wolframalpha.com/input/?i=Plot+y+Log%5Bx%5D+%3D+Log%5By%5D%3B+x%3DSqrt%5B2%5D+from+y%3D0+to+5 but we have restriction result from above. so it cannot be 4. beyond the vertical tangent in this curve ... the function diverges. http://www.wolframalpha.com/input/?i=Plot+y+Log%5Bx%5D+%3D+Log%5By%5D

OpenStudy (experimentx):

\( \sqrt 2 < e^{1/e} \implies f(\sqrt 2 ) = 2 < e < 4 \) @eliassaab prof i think this is true ... here is a short code that I wrote ... after 100 terms it shows that it's value is 2 http://ideone.com/28ZDlV also I think we can also use monotone convergence theorem to show that it's limit point exits.

OpenStudy (experimentx):

I am not so sure ... if it's the right proof ... let \( \displaystyle a_n = (\sqrt 2) ^ {a_{n-1}} \) with \( a_1 = \sqrt 2 \). let \( a_{n-1} < 2 \) , then \(a_n = (\sqrt 2) ^{a_{n-1}} < (\sqrt 2) ^{2} = 2\). Hence by induction, the sequence is bounded. for monotonicity ...

OpenStudy (experimentx):

*for monotonicity ... it's obvious. \( a_{n} > a_{n-1} \) hence it must converge.

OpenStudy (anonymous):

I just logged on to say that \[ z= \sqrt 2 \] then an infinite tower of z will converge to 2

OpenStudy (anonymous):

Try this code with Mathematica with \( z=\sqrt 2\). We can construct the following table with h[n_, z_] := Nest[z^# &, z, n] h gives a tower of power of z of length n+1 {{10, 1.98871}, {11, 1.99219}, {12, 1.99459}, {13, 1.99626}, {14, 1.99741}, {15, 1.9982}, {16, 1.99876}, {17, 1.99914}, {18, 1.9994}, {19, 1.99959}, {20, 1.99971}, {21, 1.9998}, {22, 1.99986}, {23, 1.9999}, {24, 1.99993}, {25, 1.99995}, {26, 1.99997}, {27, 1.99998}, {28, 1.99998}, {29, 1.99999}, {30, 1.99999}, {31, 1.99999}, {32, 2.}, {33, 2.}, {34, 2.}, {35, 2.}, {36, 2.}, {37, 2.}, {38, 2.}, {39, 2.}, {40, 2.}, {41, 2.}, {42, 2.}, {43, 2.}, {44, 2.}, {45, 2.}, {46, 2.}, {47, 2.}, {48, 2.}, {49, 2.}, {50, 2.}} @experimentX

OpenStudy (anonymous):

The code to generate the above code is Table[{i, h[i, N[Sqrt[2]]]}, {i, 10, 50}]

OpenStudy (anonymous):

Looking at general case:\[ \begin{array}{rcl} z^{z^{z^{\cdot ^{\cdot^{\cdot}}}}} &=& c \\ z^{\left(z^{z^{z^{\cdot ^{\cdot^{\cdot}}}}}\right)} &=& z^c \\ z^{z^{z^{\cdot ^{\cdot^{\cdot}}}}} &=& z^c \\ c &=& z^c \\ z^c-c &=& 0 \\ \end{array} \] Seems like you'd need to use Newton's method to find a solution.

OpenStudy (callisto):

Would you mind elaborating it? By the way, I haven't learnt Newton's method yet.

OpenStudy (anonymous):

It's an iterative root finding method. You guess \(x_0\), then you use:\[ x_{n+1} = x_n -\frac{f(x_n)}{f'(x_n)} \]To get closer to the root. In this case \(f(x) = x^c-c\).

OpenStudy (anonymous):

Wait nevermind what I said.

OpenStudy (anonymous):

Take both sides to the power of \(1/c\). \[ z = c^{1/c} \]

OpenStudy (anonymous):

so if you had \(=4\) then it would be \(4^{1/4}\)

OpenStudy (callisto):

Which is \(\sqrt{2}\)?

OpenStudy (unklerhaukus):

not all infinities are the same

OpenStudy (anonymous):

Hmmm, this is a tough question.

OpenStudy (anonymous):

So far, no one has elaborated on the fallacy.

OpenStudy (anonymous):

For one, is there a value for \(z\) which will even give us 4? It seems \(\sqrt{2}\) doesn't do this.

OpenStudy (callisto):

Hmm, but isn't it what we get when we solve the equation? Why is it not correct? What's wrong when we solve the equation?

OpenStudy (anonymous):

Well, everything we did when we solved to get \(z=c^{1/c}\) assumed that the tower converged.

OpenStudy (anonymous):

certain assumptions were made. I'm not sure which assumption was wrong.

OpenStudy (anonymous):

I have a question. When your teacher solved for c, did he use my method, or the log method?

OpenStudy (callisto):

Here's what he wrote: \[z^{z^{z^{z^{.^{.^{.}}}}}}=2\]\[z^2=2\]\[z=\sqrt{2}\]

OpenStudy (anonymous):

This is a rough approximation.

OpenStudy (anonymous):

When we consider the series @experimentX talked about, we could find out where it hypothetically stops being convergent.

OpenStudy (anonymous):

All I can really think to say is that the flaw with: \[z^{z^{z^{z^{.^{.^{.}}}}}}=4\]\[z^4=4\]\[z=\sqrt{4}\] lies with the assumption that \(z\) exists.

OpenStudy (anonymous):

Consider if we assume for example that and \(x\) exists such that:\[ 0x = 4 \]

OpenStudy (callisto):

Now, I am very confused. How do you define and equation? How do you know if there exist a solution to the equation? How are convergence and divergence related to solving an equation?

OpenStudy (anonymous):

convergence and divergence are related to anything that involve infinity.

OpenStudy (anonymous):

There is no way to know solutions exist.

OpenStudy (anonymous):

Unless some other math person has found solutions

OpenStudy (anonymous):

Okay think of the equation:\[ 0x = 4 \]once more. If \(x=\infty\) then we can possibly make this equation work.

OpenStudy (anonymous):

From now on let's define \[ h(z) = z^{z^{z^{z^{.^{.^{.}}}}}} \]

OpenStudy (anonymous):

Consider we tried \(\lim_{z\to -\infty} h(z)\) It would swap between 0 and an indeterminate form.

OpenStudy (anonymous):

If we tried \(\lim_{z\to 0}h(z)\) we would get an indeterminate form

OpenStudy (anonymous):

What if \(h(z)\) isn't actually a function? For example consider \(\pm x=2\). We let \(x=2\) and get \(-2\) or \(2\) and then proceed to say \(-2=2\). Yet in the case of \(x=0\), we know \(\pm x=0\) and is a function.

OpenStudy (anonymous):

Remember than indeterminate forms such as \(0/0\) are problematic because they can give you any number.

OpenStudy (anonymous):

What if the solution to the equation is a complex number? In that case, our exponent rules would not work.

OpenStudy (anonymous):

I wish I had a better answer than "the assumption z exists leads to a fallacy", because it doesn't properly pinpoint what is going wrong here.

OpenStudy (callisto):

Sorry, I am having trouble in understanding the example ±x=2.

OpenStudy (anonymous):

Well, if we said that \(g(x) = \pm x\), then we are assuming \(\pm x\) is a function, and it leads to many issues. We could say find \(z\) where \(g(z) = c\). Then we find out \(g^{-1}(c) = z = |c|\). In the case where \(c=2\), then \(z=2\), in the case where \(c=-2\) then \(z=2\). We would then say \(2=g(2)=-2\). See the problem here?

OpenStudy (callisto):

Ah, by the way, when we draw conclusion that 2=4 from \(z^{z^{z^{\cdot ^{\cdot^{\cdot}}}}} = 2\) and \(z^{z^{z^{\cdot ^{\cdot^{\cdot}}}}}=4\), we do it by solving the two equations, getting both z's being equal to \(\sqrt{2}\). So, get left side of both equations are the same, so, the right side should also be the same, i.e. 2=4.

OpenStudy (callisto):

In your case, by solving your equation, you actually get two different answers.

OpenStudy (anonymous):

Remember when you said \[ z^2=2 \]This gives 2 answers, doesn't it? \[z=\pm \sqrt{2} \]

OpenStudy (callisto):

Hmm... Yes.

OpenStudy (anonymous):

When you said \[ z^4 = 4 \]You have 4 answers: \[ z^2=\pm 2 \implies z=\pm \sqrt{\pm 2} \]

OpenStudy (anonymous):

Which is just \(\pm \sqrt{2}\) and \(\pm \sqrt {2}i\).

OpenStudy (anonymous):

\[ z^c-c=0 \implies z=c^{\frac{1}{c}} \] You do not Newton's method to do that.

OpenStudy (callisto):

@wio And then?

OpenStudy (anonymous):

@eliassaab The problem is that \(1/c\) exponent is troublesome, especially if \(c=2^n\).

OpenStudy (anonymous):

Newton's method makes it clear there will be many solutions, while that exponent ignores this fact.

OpenStudy (anonymous):

@Callisto I would like to know the actual answer, but I only have many hypothesis.

OpenStudy (anonymous):

My best answer is "you can't prove something is true by assuming it is true". You started out assuming \(h(z)=4\), and then used algebra to show \(h^{-1}(4)=\sqrt{2}\). The reality is that \(h(\sqrt{2})=2\neq 4\).

OpenStudy (callisto):

So, there must be something wrong when we solve it?

OpenStudy (anonymous):

Yeah.

OpenStudy (anonymous):

Consider \[\begin{array}{rcl} \prod\limits_{k=0}^{\infty}x&=&c\\ x\prod\limits_{k=0}^{\infty}x&=&c\\ xc&=&c\\ x&=&1 \end{array} \]Now: \[ \prod\limits_{k=0}^{\infty}x=5\implies x\prod\limits_{k=0}^{\infty}x=5\implies 5x=5\implies x=1 \]And:\[ \prod\limits_{k=0}^{\infty}x=6\implies x\prod\limits_{k=0}^{\infty}6=5\implies 6x=6\implies x=1 \] Thus: \[ 5=\prod\limits_{k=0}^{\infty}x=6 \]

OpenStudy (anonymous):

The fallacy at play here isn't any different than yours, is it?

OpenStudy (anonymous):

Doing it for sums: \[ \begin{array}{rcl} \sum\limits_0^{\infty}x&=&c\\ x+\sum\limits_0^{\infty}x&=&c\\ x+c&=&c\\ x&=&0 \end{array} \]

OpenStudy (anonymous):

So you have a problem with dividing by \(c\) because \(c\) might be \(0\), but not taking both sides to the \(1/c\) power... when \(c\) might be \(0\)?

OpenStudy (anonymous):

What If I just say \(c\neq 0\)? Does that change much?